OpenStudy (anonymous):

Jerry has a total of 23 marbles. The marbles are either blue or green. He has three more blue marbles than green marbles. How many green marbles does he have?

OpenStudy (lgbasallote):

let x = blue y = green x + y = 23 it is also said that there are three more blue than green so x = 3 + y let's sub this value of x into the first equation 3 + y +y = 23 3 + 2y = 23 2y = 23 - 3 2y = 20

OpenStudy (lgbasallote):

if you divide both sides by 2 you get y = 10 tehrefore there are 10 green marbles

OpenStudy (anonymous):

okay i think it's like this suppose there are x green marbles and since there are three more blue marbles than green marbles, we can say that he has x+3 blue marbles so basically green=x blue=x+3 and total is 23 so x+x+3=23 so x=10 10 green marbles